each of the congruent angles of an isosceles triangle measures 9 less than 4 times the vertex angle. Find the measures of the angles of the triangle.
Answer by rapaljer(4671)

Let x = measure of the vertex angle
4x-9 = measure of each congruent angle

Equation = Sum of the three angles = 180
x + 4x -9 + 4x - 9 = 180
9x -18=180
9x -18 +18 = 180+18
9x = 198

Divide both sides by 9:
x = 22 degrees
4x-9 =4*22 - 9 = 88-9 = 79 degrees

Check: 22+79+79 = 180
It checks!

R^2 at SCC